OpenStudy (anonymous):

(x^1/2)^1/2)^1/2=x^7/8 how?

OpenStudy (anonymous):

x=1 .....ans

OpenStudy (anonymous):

(x^1/2)^1/2)^1/2=x^7/8 ..squaring both sides 3times to get x=x^(((7/8)2)2)2 x=x^7 x^7-x=0 x=1 root x=0 root x=-1 root the other 3 roots are complex numbers

OpenStudy (bahrom7893):